What’s the interest rate of a 8-year, annual $3,600 annuity with present value of $20,000? (Use a time value of money calculator or a spreadsheet. Round your answer to 2 decimal places.)
Answer #1
we need to use present value of annuity formula
= P [1-(1+r)^-n]/r
Where,
r= interest rate per period
P= periodic payment i.e. $36000
n= number of period i.e. 8 Years
$ 20000 = 3600[1-(1+r)^-8]/r
$20000/3600 = [1-(1+r)^-8]/r
5.5556* r = 1-(1+r)^-8
r= 8.90%
Therefore inerest rate is 8.90%